Solve (D^3-6D^2D'+12DD'^2-8D'3)z=e^2x+y

1
Expert's answer
2022-02-07T17:45:21-0500

Given, "(D^3-6D^2D'+12DD'^2-8D'^3)z=e^{2x+y}"

The auxiliary equation is "m^3-6m^2+12m-8=0"

The roots of the equation are m = 2 (thrice)

Therefore, the complementary function is

"C.F = f_1(y+2x)+xf_2(y+2x)+x^2f_{3}(y+2x)"

Particular integral

"\begin{aligned}

P.I &= \dfrac{1}{D^3-6D^2D’+12DD’^2-8D’^3} e^{2x+y}\\

&= \dfrac{1}{2^3-6\cdot 2^2\cdot 1+12\cdot 2 \cdot 1^2-8\cdot 1^3} e^{2x+y}\\
&\text{(Replacing D = 2, D' = 1)}\\
&= \dfrac{1}{0} e^{2x+y}~~ \text{(Rule fails)}\\
&= x \dfrac{1}{3D^2 – 12DD’ +12D’^2} e^{2x+y} \\
&\text{(Differentiating with respect to D and multiplying by}~ x)\\
&= x\dfrac{1}{3 \cdot 2^2-12 \cdot 2 \cdot 1 + 12 \cdot 1^2} e^{2x+y}\\
&~~~~~~~\text{(Replacing $D = 2, D'= 1$)}\\
&= x~\frac{1}{0} e^{2x+y}~~ \text{(Rule fails)}\\
&= x^2 \dfrac{1}{6D – 12D'} e^{2x+y}\\ &\text{(Differentiating with respect to D and multiplying by $x$)}\\
&= x^2 \dfrac{1}{6 \cdot 2 – 12 \cdot 1} e^{2x+y} ~~\text{(Rule fails)}\\
&= x^3~ \dfrac{1}{6} e^{2x+y} \\
&\text{(Differentiating with respect to D and multiplying by $x$)}\\
\therefore ~P.I & = \dfrac{x^{3}}{6}e^{2x+y}
\end{aligned}"


Hence the solution is given by,

"\begin{aligned}
z &= C.F + P.I\\
&= f_1(y+2x)+xf_2(y+2x)+x^2f_{3}(y+2x) + \dfrac{x^{3}}{6}e^{2x+y}
\end{aligned}"